OpenStudy (anonymous):

Need help for Calc 1 quiz!!! Implicit differentiation of y=tan^-1(x) / x^2 I did the product rule, but I'm lost in how to simplify it....

OpenStudy (mathwizzard3):

https://www.wolframalpha.com/

OpenStudy (mathwizzard3):

That will help with calculations like that! :3

OpenStudy (anonymous):

Yes, but it doesn't show steps.

OpenStudy (anonymous):

y = f.g \[f = \tan^{-1} x \\ g = x^{-2}\] apply product rule, \[y \prime = f \prime g + g \prime f \\ f \prime = \frac{1}{1 + x^2} \\ g \prime = -2.x^{-3}\] Now can you simplify?

OpenStudy (jdoe0001):

\(\bf y=\cfrac{tan^{-1}(x)}{x^2} \\ \quad \\ \cfrac{dy}{dx}=\cfrac{dy}{dx}\left[ \cfrac{tan^{-1}(x)}{x^2} \right]\implies \cfrac{dy}{dx}=\cfrac{dy}{dx}[tan^{-1}(x)\cdot x^{-2}] \\ \quad \\ \left( \cfrac{1}{1+x^2}\cdot x^{-2} \right)+\left( tan^{-1}(x)\cdot -2x^{-3}\right) \\ \quad \\ \cfrac{1}{x^2(1+x^2)}+tan^{-1}(x)\cdot -2\cdot \cfrac{1}{x^3}\implies \cfrac{1}{x^2+x^4}-\cfrac{2tan^{-1}(x)}{x^3}\)

OpenStudy (anonymous):

Thank you jdoe0001 !! The way you did it was wayyy better to understand how parts fit together. :)

OpenStudy (anonymous):

I did good too, that's not fair. :)